HETAV PATEL

Grade 11,

The distance covered by a particle in t seconds is given by x = 3 + 8t – 4t2. After 1 second velocity will be (A) 0 unit/second (B) 3 units/second (C) 4 units/second (D) 7 units/second

The distance covered by a particle in t seconds is given by x = 3 + 8t – 4t2. After 1 second velocity will be
(A) 0 unit/second (B) 3 units/second (C) 4 units/second (D) 7 units/second

Grade:

2 Answers

Saurabh Koranglekar
3 years ago
Givenx=3+8t−4t2x=3+8t-4t2,
On differentiating both sides w.r.tttwe get

Att=1,v=8−8−0t=1,v=8-8=0unit.
Vikas TU
3 years ago
Der student 
Correct ans is option A …........................................................................................................